Find the mean for this list of numbers 51 90 42 38 43 61 6 36 66 55 48 11 84 29 46 Mean -
Find the median for this list of numbers 9 28 78 81 41 88 99 27 89 24 30 39 2 63 91 Median-
Answer #1

1.

Mean=Sum of terms / Number of terms

Sum of terms = 51 + 42 + 43 + 6 + 66 + ⋯+ 29 = 706

Number of terms=15

Mean=Sum of terms / Number of terms

= 706 / 15

= 47.0667

2.

The median is the middle number in a sorted list of numbers.

So, to find the median, first place the numbers in value order and then find the middle number.

Ordering the data from least to greatest, we get:

2   9   24   27   28   30   39   41   63   78   81   88   89   91   99   

So, the median is 41 .